OpenStudy (nannie):

I am suppose to simplify and type it as an exponential notation with positive exponents. a^14 •a^0 = 0 I believe my answer should be 0.000000e+0• 0.000000e+0= 0.000000e+0

OpenStudy (unklerhaukus):

what is •

OpenStudy (nannie):

?

OpenStudy (unklerhaukus):

your question has a symbol i am unfamiliar with, what does '•' indicate in your question. does it mean 'times' or something else

OpenStudy (nannie):

*

OpenStudy (unklerhaukus):

is that the symbol for taking the complex conjugate

OpenStudy (nannie):

It is to multiply. Times I was going to answer a^14=0 but that is not an exponential notation with positive exponents.

OpenStudy (unklerhaukus):

a^0 = 1 (unless a is 0 in which case 0^0 is indeterminate)

OpenStudy (unklerhaukus):

a^14=0 a = 0

OpenStudy (unklerhaukus):

a =0•10^0

OpenStudy (nannie):

It shows a^14xa^0=? Then type it as an exponential notation with positive exponents

OpenStudy (unklerhaukus):

0^1

OpenStudy (nannie):

it is a Math Lab question.

OpenStudy (nannie):

ok.that makes sense. THANKS